(b) A block of ice of mass 40g at 00C is placed in a calorimeter containing 400g of water at 200C.
Ignoring the heat absorbed by the calorimeter, determine the final temperature of the mixture after all the ice have melted. (Specific latent heat capacity of fusion of ice = 340,000JKg-1 and the specific heat capacity of water =4200 J/KKg (6mks)
1
Expert's answer
2018-12-19T10:30:11-0500

The heat balance equation

λm+cm(t_3-t_1 )=cM(t_2-t_3)

Thus

t_3=(cMt_2+cmt_1-λm)/(cm+cM)

=(4200(0.4×20+0.04×0)-340000×0.04)/(4200(0.04+0.4))=11℃